# Fold catastrophe in breaking waves

### v1

@@ Abstract
We present a dynamical-systems perspective on wave breaking for ideal incompressible free-surface flows. By tracking the most energetic hotspot on the wave surface, we find that near breaking the surface slope $m$ evolves on a fast timescale governed by the small parameter $\epsilon=(\partial_{z}u)^{-1}$, the inverse vertical velocity gradient at the hotspot, while the focusing parameter $A=(U-C_{e})/(U-C_{req})$ varies slowly and adiabatically. Here $U$ is the horizontal fluid velocity at the energetic point, $C_{e}$ its propagation speed, and $C_{req}$ the equivalent crest speed. This slow–fast structure reveals a fold catastrophe in the $(m,A)$ space whose boundary forms the geometric skeleton organizing the dynamics near breaking. Finite-time blowup occurs when the trajectory crosses this boundary, marking the onset of breaking.
@@ 1 Introduction
Breaking onset occurs when the fluid velocity at the crest equals the crest speed, causing fluid particles to detach from the surface  (phillips1966dynamics; PhillipsBanner1974; BannerPhillips1974). When a wave breaks on a sloping beach, the fluid may plunge forward enclosing an air bubble or it may spill. In deep water, the spilling motion is typical, with a characteristic formation of a bulge and fluid elements sliding down the leading slope  (duncan1981Royalexperimental; Duncanqiao2001gentlebreaker_crestflow; duncan1999gentlebreaker_crest_profile). At larger scales of breaking, air entrainment may take place and the breaking wave is clearly visible as a whitecap.
@@
Recent studies have uncovered fundamental properties of wave breaking by examining the dynamics of wave groups by tracking specific moving points on their surface, such as crests and energetic hotspots. Banner2014CrestSlowdown, through numerical simulations and field observations, showed that wave crests approaching their maximum height slow down significantly and either break at this reduced speed or accelerate forward unbroken. This crest slowdown is a generic feature that arises from the natural dispersion of waves (FedeleEPL2014; FedeleJFMcrestspeed2020). It explains why breaking wave crests travel about $20\%$ slower than predicted by linear theory (RappMelville1990). A following study by BarthelemyJFM2018 identified a robust precursor to the breaking onset of wave crests. This is an inception point marking a state of no return beyond which the crest inevitably approaches the breaking onset and collapses. Through detailed numerical wave tank simulations, they found a robust threshold parameter for this inception valid for wave packets propagating either in deep or intermediate water depths. BarthelemyJFM2018 found that maximally tall non-breaking waves are clearly separated from marginally breaking waves when the parameter $B=U/C_{r}\approx 0.85$, where $U$ is the horizontal fluid velocity and $C_{r}$ is the crest speed of the tallest wave (see, also SaketJFM2017 for an experimental validation). DerakhtiJGR2020 validated the proposed breaking threshold for surface gravity water waves in arbitrary water depth, including shallow water breaking over varying bathymetry. Touboul_Banner_2021 proved that it also holds for waves propagating in the presence of a constant background vorticity. Derakhti2018 also showed that the breaking strength of 22D and 33D gravity waves, measured by the energy dissipated per breaking event, is strongly correlated with the rate of change $dB/dt$ at the breaking onset. Recent experimental work has examined the kinematic aspect of the breaking onset (Shemer2014LagrangianKinematics; Liberzon2017), leading to novel detection methods (Liberzon_breaking2019) that identify breaking crest bulges by analyzing instantaneous frequency variations in the wave signal (Zimmermann_Seymour2002).
@@
More recently, BoettgerJFM2023 identified a robust energetic signature of breaking inception by numerically solving for nonlinear wave packets under various forcing conditions, including cases with surface tension. They found an energetic hotspot on the forward face of the crest where the kinetic energy density attains its maximum (Boettger2024). Crest breaking invariably occurred when the kinetic energy growth rate, measured at this hotspot, exceeded a critical threshold. This threshold, apparently dependent on the wave generation method, reflects an excessive convergence of kinetic energy at the crest that a wave cannot sustain. BoettgerJFM2023; BoettgerPRF2024 analyzed the geometric and kinematic properties of this energetic inception point finding a weakly dependence from how waves are generated or forced.
@@
In contrast, McAllister2023 performed numerical simulations of 22D inviscid, and irrotational water waves in the absence of surface tension. By tracking the surface point of maximum inflection, they found that the minimum local interface slope $m_{b}=\tan(\pi/6)\approx 0.5774$ at that point separates breaking from non-breaking waves. This threshold corresponds to an angle of $30^{\circ}$, matching the limiting steepness derived by Stokes1847. However, this threshold is not universal, as it strongly depends on how the waves are forced (duncan1983JFMbreaking). Indeed, Boettger2025 noted that the critical slope $m_{b}$ varies in the presence of surface tension and close to an angle of $60^{\circ}$. They showed that surface tension effects are significant as the wave crest approaches its maximum steepness. For the steepest non-breaking waves, this manifests as a bulge on the crest tip that increases the local interface angle.
@@
A theoretical effort to explain the wave breaking phenomenon is related to the superharmonic instability of periodic Stokes waves (LonguetHigginsSuperHarm1978; Creedon_Deconinck_Trichtchenko_2022; DeconinckStokes2023; berti2024infinitely). Recent studies have investigated the emergence of breaking water waves in deep water induced by perturbations to unstable Stokes waves (MansarBridges2025SHInstability; DeconinckStokes2023), originally studied by Cokelet1978. In this regard, FedeleJFM2014 considered the dynamics of the compact Dyachenko-Zakharov (cDZ) equation for unidirectional deep-water waves, a third-order model derived from the Euler equations (cDZ2011). Perturbations to superharmonically unstable uniform wave trains of the cDZ equation in deep water evolve according to a modified Korteweg–de Vries equation, suggesting a possible primordial mechanism conducive to wave breaking, inherited from the Euler equations (FedeleJFM2014).
@@
Another complementary approach has been the study of breaking in wave systems governed by nonlinear dispersive equations. In this context, a wave breaks when its profile steepens until the slope becomes infinite in finite time, leading to the blow-up of smooth solutions. This behavior is of significant interest in fluid dynamics, particularly for the Euler equations modeling surface water waves (see, e.g., Dyachenko2008), as well as asymptotic models such as the Whitham1967 equation among others (KalischWhitham2015). Seliger1968 showed that if the initial wave profile is sufficiently asymmetric, it leads to breaking in the Whitham equation, with nonlinear steepening overcoming dispersion. Seliger’s method focuses on tracking the maximum slopes at the inflection points on the forward and backward faces of a wave crest, where the curvature vanishes. He derived differential equations governing the evolution of the slope difference and showed this quantity becomes unbounded in finite time. Constantin1998 and Hur2017 later provided rigorous proofs of this result. Measuring a surface slope, or angle, to detect the breaking inception, or point of no return, of a wave is appealing. It is relatively simple compared to the $B$ threshold, which requires measuring both fluid and crest velocities. However, whether a universal slope threshold exists that separates non-breaking from breaking waves remains to be established.
@@
In this work, we aim to provide further insight into the nature of wave breaking and the existence of thresholds for breaking inception. We adopt a dynamical-systems perspective and, from first principles, derive a slow–fast model that captures wave dynamics near breaking, as observed from the energetic hotspot (BoettgerJFM2023).
@@
To derive this model, we first review the Euler equations for 22D irrotational surface water waves. We then consider the evolution of local observables following a generic moving point on the surface. These observables evolve according to a system of coupled ordinary differential equations, which is generally not closed. We then provide the physical insights that led to the identification of a relational fast timescale $\tau$ near wave breaking (Rovelli1991TimeHypothesis; rovelli2004quantum; rovelli2018order). This defines a steepening time that is faster than the absolute Newtonian time. The wave evolution on the timescale $\tau$, as observed at the energetic hotspot, is characterized by slow variations of kinematic variables and rapid changes of geometric parameters such as the surface slope. A slow–fast model naturally arises from first principles, revealing breaking as a fold catastrophe (Thom1975; zeeman1976catastrophe; arnold1984catastrophe; Gilmore1993).
@@
Finally, we discuss our results in relation to the numerical findings of BoettgerJFM2023 and McAllister2023, and outline how wave breaking can be viewed as the gravity analogue of collapsing black holes.
@@ 2 Governing Equations
where $L,T$ are a characteristic wavelenghth and wave period, and the speed $c=L/T$. The Euler equations remain invariant under this rescaling, except that the gravitational acceleration transforms as $g\to gT/c$. Following BarthelemyJFM2018, we choose $L=2\pi/k$ and $T=2\pi/\omega$, where the wave frequency $\omega$ and wavenumber $k$ satisfy the linear dispersion relation $k=\omega^{2}/g$ in deep water. Then, $c=\omega/k=\sqrt{gk}$ is the phase speed and $g=2\pi$.
@@ 2.1 Kinetic energy transport
where $k_{e}=|\mathbf{u}|^{2}/2=(u^{2}+w^{2})/2$ is the kinetic energy density per unit mass. The kinetic energy transport equation (Tulin2007; FedeleJFM2016)
@@ 2.2 Pressure relations
$\eta(x,:)\mapsto\eta(-x,:)\,,\,\,\{u(x,:)\mapsto-u(-x,:),\,\,w(x,:)\mapsto w(-x,:)\},\,\,p(x,:)\mapsto p(-x,:),$
@@ 2.3 Vorticity at the Free Surface
For an inviscid and irrotational fluid (water), vorticity in the bulk vanishes identically by Kelvin’s circulation theorem, and Eq. (5) holds everywhere in the interior of the fluid. However, the zero-stress condition at the surface allows fluid parcels to rotate with the evolving unit normal vector (LonguetHiggins1998), thereby generating surface vorticity (FedeleJFM2016):
@@
The free surface also supports a vortex sheet with strength $\gamma=\mathbf{u}_{t}^{+}-\mathbf{u}_{t}^{-}$, defined as the jump in tangential velocity across the interface (Terrington2020; Brons2014; lundgren1999generation). Here, $\mathbf{u}_{t}^{+}$ and $\mathbf{u}_{t}^{-}$ denote the tangential velocity just above and just below the surface, respectively. Because the density and viscosity of the upper fluid (air) are effectively negligible, the sheet strength reduces to $\gamma=-\mathbf{u}_{t}^{-}$, that is, minus the tangential velocity of the water at the surface. This concentrated vortex layer formally contains all the vorticity generated by surface curvature and by the unsteady deformation of the free boundary.
In a viscous fluid, surface-generated vorticity diffuses into the interior, producing shear layers (phillips1966dynamics; fedele2022momentary). These can grow and interact with the vortex sheet while conserving the total circulation induced both by the vortex sheet and by the viscous interior flow (Terrington2020; Brons2014; lundgren1999generation).
@@ 3 Tracking observables on the free surface to probe breaking
We study the surface dynamics by tracking the values of physical fields, or observables, in time at a specific moving point $Q$ on the wave surface. This defines a trajectory that is not Lagrangian, as it follows a geometrically distinguished point on the wave profile rather than a fluid particle. For example, BarthelemyJFM2018 investigated breaking inception by tracking a wave crest. Seliger1968 and McAllister2023 tracked surface points of maximum inflection, and BoettgerJFM2023 followed surface points of maximum kinetic energy.
@@
The time evolution of these quantities leads to a system of ordinary differential equations, which is not closed (see appendix 13). The time evolution of each of these local observables depend on the others due to the nonlinear nature of the Euler equations. Completing the ODE system would require adding dynamical equations for all these related observables, which is an extremely cumbersome task. A practical solution would involve a truncation procedure to achieve closure. Nevertheless, near the breaking regime, a slow-fast timescale separation occurs and the time evolution of a few key observables is sufficient to characterize wave breaking as a cusp catastrophe bifurcation.
@@ 3.1 Dimensionless wave parameters
The ratio $s$ is kinematic and measures the relative importance of vertical fluid velocities with respect to the horizontal fluid flow. The slope of the fluid velocity vector $\vec{\mathbf{u}}$ is given by $\tan\theta_{\vec{\mathbf{u}}}=s$, where $\theta_{\vec{\mathbf{u}}}$ is the angle with respect to the horizontal. The ratio $m_{2}$ is a momentum parameter
@@
with $C_{req}$ being the equivalent crest speed at $Q$, i.e., the speed as if there were a crest at $Q$, and derived in Appendix 15. In general, near a focusing wave crest, $C_{req}\geq C$ at $Q$, and when $Q$ passes through the wave crest, $C=C_{req}$ and $A=1$ where the maximum crest height is attained at focus. [^1] From the definition of $A$, $C=(1-A)U+AC_{req}$, meaning $C$ is a weighted average between the fluid speed $U$ and the equivalent crest speed $C_{req}$. In the range of $0<A<1$, we have $U<C<C_{r}$ and $U/C<1$ and $C_{req}/C>1$. At the breaking onset, where $U\rightarrow C$, $A\rightarrow 1$.

### v2

@@ Abstract
We present a dynamical-systems perspective on wave breaking for ideal incompressible free-surface flows. By tracking the most energetic hotspot on the wave surface, we find that near breaking the surface slope $m$ evolves on a fast timescale governed by the small parameter $\epsilon=(\partial_{z}u)^{-1}$, the inverse vertical velocity gradient at the hotspot, while the focusing parameter $A=(U-C_{e})/(U-C_{req})$ varies slowly and adiabatically. Here, $U$ is the horizontal fluid velocity at the energetic point, $C_{e}$ its propagation speed, and $C_{req}$ the equivalent crest speed. This slow–fast structure reveals a fold catastrophe in the $(m,A)$ space, whose boundary forms the geometric skeleton organizing the dynamics near breaking. Finite-time blowup occurs when the trajectory crosses this boundary, marking the onset of breaking.
@@
2020 MSC:
37E35 ,
2020 MSC:
76B07
2020 MSC:
76B15 ,
2020 MSC:
58K35
journal: Communications in Nonlinear Science and Numerical Simulation
[…]
organization=School of Civil and Environmental Engineering, Georgia Institute of Technology,addressline=790 Atlantic Drive, city=Atlanta, postcode=30332-0355, state=Georgia, country=USA
@@ 1 Introduction
Breaking onset occurs when the fluid velocity at the crest just exceeds the crest speed, often causing fluid particles to detach subsequently from the surface  (Phillips, 1966; Phillips and Banner, 1974; Banner and Phillips, 1974). When a wave breaks on a sloping beach, the fluid may plunge forward enclosing an air bubble or it may spill. In deep water, the spilling motion is typical, with a characteristic formation of a bulge and fluid elements sliding down the leading slope  (Duncan, 1981; Qiao and Duncan, 2001; Duncan et al., 1999). At larger scales of breaking, air entrainment may take place and the breaking wave is clearly visible as a whitecap.
@@
Recent studies have uncovered fundamental properties of wave breaking by examining the dynamics of wave groups by tracking specific moving points on their surface, such as crests and energetic hotspots. Banner et al. (2014), through numerical simulations and field observations, showed that wave crests approaching their maximum height slow down significantly and either break at this reduced speed or accelerate forward unbroken. This crest slowdown is a generic feature that arises from the natural dispersion of waves (Fedele, 2014a; Fedele et al., 2020). It explains why breaking wave crests travel about $20\%$ slower than predicted by linear theory (Rapp and Melville, 1990). A following study by Barthelemy et al. (2018) identified a robust precursor to the breaking onset of wave crests. This is an inception point marking a state of no return beyond which the crest inevitably approaches the breaking onset and collapses. Through detailed numerical wave tank simulations, they found a robust threshold parameter for this inception valid for wave packets propagating either in deep or intermediate water depths. Barthelemy et al. (2018) found that maximally tall non-breaking waves are clearly separated from marginally breaking waves when the parameter $B=U/C_{r}\approx 0.85$, where $U$ is the horizontal fluid velocity and $C_{r}$ is the crest speed of the tallest wave (see, also Saket et al. (2017) for an experimental validation). Derakhti et al. (2020) validated the proposed breaking threshold for surface gravity water waves in arbitrary water depth, including shallow water breaking over varying bathymetry. Touboul and Banner (2021) proved that it also holds for waves propagating in the presence of a constant background vorticity. Derakhti et al. (2018) also showed that the breaking strength of 22D and 33D gravity waves, measured by the energy dissipated per breaking event, is strongly correlated with the rate of change $dB/dt$ at the breaking onset. Recent experimental work has examined the kinematic aspect of the breaking onset (Shemer and Liberzon, 2014; Itay and Liberzon, 2017), leading to novel detection methods (Liberzon et al., 2019) that identify breaking crest bulges by analyzing instantaneous frequency variations in the wave signal (Zimmermann and Seymour, 2002).
@@
More recently, Boettger et al. (2023) identified a robust energetic signature of breaking inception by numerically solving for nonlinear wave packets under various forcing conditions, including cases with surface tension. They found an energetic hotspot on the forward face of the crest where the kinetic energy density attains its maximum (Boettger, 2024). Crest breaking invariably occurred when the kinetic energy growth rate, measured at this hotspot, exceeded a critical threshold. This threshold, apparently dependent on the wave generation method, reflects an excessive convergence of kinetic energy at the crest that a wave cannot sustain. Boettger et al. (2023, 2024) analyzed the geometric and kinematic properties of this energetic inception point finding a weak dependence on how waves are generated or forced.
@@
In contrast, McAllister et al. (2023) performed numerical simulations of 22D inviscid, and irrotational water waves in the absence of surface tension. By tracking the surface point of maximum inflection, they found that the minimum local interface slope $m_{b}=\tan(\pi/6)\approx 0.5774$ at that point separates breaking from non-breaking waves. This threshold corresponds to an angle of $30^{\circ}$, matching the limiting steepness derived by Stokes (1847). However, this threshold is not universal, as it strongly depends on how the waves are forced (Duncan, 1983). Indeed, Boettger et al. (2025) noted that the critical slope $m_{b}$ varies in the presence of surface tension and can reach angles approaching $60^{\circ}$. They showed that surface tension effects are significant as the wave crest approaches its maximum steepness. For the steepest non-breaking waves, this manifests as a bulge on the crest tip that increases the local interface angle.
@@
A theoretical effort to explain the wave breaking phenomenon is related to the superharmonic instability of periodic Stokes waves (Longuet-Higgins, 1978; Creedon et al., 2022; Deconinck et al., 2023; Berti et al., 2024). Recent studies have investigated the emergence of breaking water waves in deep water induced by perturbations to unstable Stokes waves (Mansar et al., 2025; Deconinck et al., 2023), originally studied by Longuet-Higgins and Cokelet (1978). In this regard, Fedele (2014b) considered the dynamics of the compact Dyachenko-Zakharov (cDZ) equation for unidirectional deep-water waves, a third-order model derived from the Euler equations (Dyachenko and Zakharov, 2011). Perturbations to superharmonically unstable uniform wave trains of the cDZ equation in deep water evolve according to a modified Korteweg–de Vries equation, suggesting a possible primordial mechanism conducive to wave breaking, inherited from the Euler equations (Fedele, 2014b).
@@
Another complementary approach has been the study of breaking in wave systems governed by nonlinear dispersive equations. In this context, a wave breaks when its profile steepens until the slope becomes infinite in finite time, leading to the blow-up of smooth solutions. This behavior is of significant interest in fluid dynamics, particularly for the Euler equations modeling surface water waves (see, e.g., Dyachenko and Zakharov (2008)), as well as asymptotic models such as the Whitham (1967) equation among others (Moldabayev et al., 2015). Seliger (1968) showed that if the initial wave profile is sufficiently asymmetric, it leads to breaking in the Whitham equation, with nonlinear steepening overcoming dispersion. Seliger’s method focuses on tracking the maximum slopes at the inflection points on the forward and backward faces of a wave crest, where the curvature vanishes. He derived differential equations governing the evolution of the slope difference and showed this quantity becomes unbounded in finite time. Constantin and Escher (1998) and Hur (2017) later provided rigorous proofs of this result. Measuring a surface slope, or angle, to detect the breaking inception, or point of no return, of a wave is appealing. It is relatively simple compared to the $B$ threshold, which requires measuring both fluid and crest velocities. However, whether a universal slope threshold exists that separates non-breaking from breaking waves remains to be established.
@@
In this work, we aim to provide further insight into the nature of wave breaking and the existence of thresholds for breaking inception. We adopt a dynamical-systems perspective and, from first principles, derive a slow–fast model that captures wave dynamics near breaking, as observed from the energetic hotspot (Boettger et al., 2023).
@@
To derive this model, we first review the Euler equations for 22D irrotational surface water waves. We then consider the evolution of local observables following a generic moving point on the surface. These observables evolve according to a system of coupled ordinary differential equations, which is generally not closed. We then provide the physical insights that led to the identification of a relational fast timescale $\tau$ near wave breaking (Rovelli, 1991, 2004, 2018). This defines a steepening time that is faster than the absolute Newtonian time $t$ measured in the laboratory frame to observe waves. The wave evolution on the fast timescale $\tau$, as observed at the energetic hotspot, is characterized by slow variations of kinematic variables and rapid changes of geometric parameters such as the surface slope. A slow–fast model naturally arises from first principles, revealing breaking as a fold catastrophe (Thom, 1975; Zeeman, 1976; Arnold, 1992; Gilmore, 1993).
@@
Finally, we discuss our results in relation to the numerical findings of Boettger et al. (2023) and McAllister et al. (2023), and outline how wave breaking can be viewed as the gravity analogue of collapsing black holes.
@@ 2 Governing Equations
where $L,T$ are a characteristic wavelenghth and wave period, and the speed $c=L/T$. The Euler equations remain invariant under this rescaling, except that the gravitational acceleration transforms as $g\to gT/c$. Following Barthelemy et al. (2018), we choose $L=2\pi/k$ and $T=2\pi/\omega$, where the wave frequency $\omega$ and wavenumber $k$ satisfy the linear dispersion relation $k=\omega^{2}/g$ in deep water. Then, $c=\omega/k=\sqrt{gk}$ is the phase speed and $g=2\pi$.
@@ 2.1 Kinetic energy transport
where $k_{e}=|\mathbf{u}|^{2}/2=(u^{2}+w^{2})/2$ is the kinetic energy density per unit mass. The kinetic energy transport equation (Tulin, 2007; Fedele et al., 2016)
@@ 2.2 Pressure relations
$\eta(x,:)\mapsto\eta(-x,:)\,,\,\,u(x,:)\mapsto-u(-x,:),\,\,w(x,:)\mapsto w(-x,:),\,\,p(x,:)\mapsto p(-x,:),$
@@ 2.3 Vorticity at the Free Surface
For an inviscid and irrotational fluid (water), vorticity in the bulk vanishes identically by Kelvin’s circulation theorem, and Eq. (5) holds everywhere in the interior of the fluid. However, the zero-stress condition at the surface allows fluid parcels to rotate with the evolving unit normal vector (Longuet-Higgins, 1998), thereby generating surface vorticity (Fedele et al., 2016):
@@
The free surface also supports a vortex sheet with strength $\gamma=\mathbf{u}_{t}^{+}-\mathbf{u}_{t}^{-}$, defined as the jump in tangential velocity across the interface (Terrington et al., 2020; Brøns et al., 2014; Lundgren and Koumoutsakos, 1999). Here, $\mathbf{u}_{t}^{+}$ and $\mathbf{u}_{t}^{-}$ denote the tangential velocity just above and just below the surface, respectively. Because the density and viscosity of the upper fluid (air) are effectively negligible, the sheet strength reduces to $\gamma=-\mathbf{u}_{t}^{-}$, that is, minus the tangential velocity of the water at the surface. This concentrated vortex layer formally contains all the vorticity generated by surface curvature and by the unsteady deformation of the free boundary.
In a viscous fluid, surface-generated vorticity diffuses into the interior, producing shear layers (Phillips, 1966; Fedele, 2022). These can grow and interact with the vortex sheet while conserving the total circulation induced both by the vortex sheet and by the viscous interior flow (Terrington et al., 2020; Brøns et al., 2014; Lundgren and Koumoutsakos, 1999).
@@ 3 Tracking observables on the free surface to probe breaking
We study the surface dynamics by tracking the values of physical fields, or observables, in time at a specific moving point $Q$ on the wave surface. This defines a trajectory that is not Lagrangian, as it follows a geometrically distinguished point on the wave profile rather than a fluid particle. For example, Barthelemy et al. (2018) investigated breaking inception by tracking a wave crest. Seliger (1968) and McAllister et al. (2023) tracked surface points of maximum inflection, and Boettger et al. (2023) followed surface points of maximum kinetic energy.
@@
The time evolution of these quantities leads to a system of ordinary differential equations, which is not closed (see B). The time evolution of each of these local observables depend on the others due to the nonlinear nature of the Euler equations. Completing the ODE system would require adding dynamical equations for all these related observables, which is an extremely cumbersome task. A practical solution would involve a truncation procedure to achieve closure. Nevertheless, near the breaking regime, a slow-fast timescale separation occurs and the time evolution of a few key observables is sufficient to characterize wave breaking as a fold catastrophe bifurcation.
@@ 3.1 Dimensionless wave parameters
The ratio $s$ is kinematic and measures the relative importance of vertical fluid velocities with respect to the horizontal fluid flow. The slope of the fluid velocity vector $\mathbf{u}$ is given by $\tan\theta_{\mathbf{u}}=s$, where $\theta_{\mathbf{u}}$ is the angle with respect to the horizontal. The ratio $m_{2}$ is a momentum parameter
@@
with $C_{req}$ being the equivalent crest speed at $Q$, i.e., the speed as if there were a crest at $Q$, and derived in D. In general, near a focusing wave crest, $C_{req}\geq C$ at $Q$, and when $Q$ passes through the wave crest, $C=C_{req}$ and $A=1$ where the maximum crest height is attained at focus. [^1] From the definition of $A$, $C=(1-A)U+AC_{req}$, meaning $C$ is a weighted average between the fluid speed $U$ and the equivalent crest speed $C_{req}$. In the range of $0<A<1$, we have $U<C<C_{r}$ and $U/C<1$ and $C_{req}/C>1$. At the breaking onset, where $U\rightarrow C$, $A\rightarrow 1$.